student tickets cost $5 and adult tickets costs $8. one night they had a total of 104 people attend the show. If they raised $583, how many tickets of each were sold
Answer by Cromlix(4381)

Hi there,
Students + Adults = 104
x + y = 104........(1)
Students tickets + Adults tickets = $583
5x + 8y = 583......(2)
.............
x + y = 104........(1)
5x + 8y = 583......(2)
Multiply Eq(1) by 8
8x + 8y = 832......(1)
5x + 8y = 583......(2)
Subtract Eq(2) from Eq(1)
3x = 249
x = 83
Substitute x = 83 into
Eq(1)
x + y = 104.........(1)
83 + y = 104
y = 104 - 83
y = 21
They sold 83 student's tickets
and 21 adult's tickets.
Hope this helps :-)